# A Concise Lyapunov Analysis of Nesterov’s Accelerated Gradient Method

### v1

@@ 2 Preliminaries
The following well-known inequalities for $L$-smooth and convex functions will be used later.
@@
$f(y)$ $\leq f(x)+(-\alpha+\frac{L\alpha^{2}}{2})\left|\nabla f(x)\right|^{2}$ (3)
which leads to
@@
when $\alpha=\frac{1}{L}$, and $f(y)\leq f(x)$ for all $\alpha\in[0,\frac{1}{2L}]$.
Lemma 5.
[8, Theorem 2.1.5, eq. (2.1.7)] If $f:\,\mathbb{R}^{n}\rightarrow\mathbb{R}$ is $L$-smooth and convex, then
$f(y)\geq f(x)+(y-x)\cdot\nabla f(x)+\frac{1}{2L}\left|\nabla f(x)\right|^{2}$
for all $x,y\in\mathbb{R}^{n}$.
@@ 3 General convex functions
Let $p_{0}=0$. Equation (7) also holds for $k=0$ by the initial condition $x_{0}=x_{-1}$ and $y_{0}=x_{0}$. By (7) and (5), it is straightforward to verify that
@@
$\,-2\alpha a_{k+1}(y_{k}-y^{*})\cdot\nabla f(y_{k})+a_{k+1}^{2}\alpha^{2}_{k}\left|\nabla f(y_{k})\right|^{2}.$ (8)
By convexity and $L$-smoothness of $f$ (Lemma 5), we have
$(y_{k}-x_{k})\cdot\nabla f(y_{k})\geq f(y_{k})-f(x_{k})+\frac{1}{2L}\left|\nabla f(y_{k})\right|^{2},$ (9)
@@
$(y_{k}-y^{*})\cdot\nabla f(y_{k})\geq f(y_{k})-f^{*}+\frac{1}{2L}\left|\nabla f(y_{k})\right|^{2}.$ (10)
@@
$\quad-\frac{2\alpha a_{k+1}(a_{k+1}-1)}{2L}\left|\nabla f(y_{k})\right|^{2}-2\alpha a_{k+1}(f(y_{k})-f^{*})$
$\quad-\frac{2\alpha a_{k+1}}{2L}\left|\nabla f(y_{k})\right|^{2}+a_{k+1}^{2}\alpha^{2}_{k}\left|\nabla f(y_{k})\right|^{2}$
@@
where the first inequality follows from substituting (9) and (10) into (8) and the second inequality used $\alpha\leq\frac{1}{L}$, $f(x_{k+1})\leq f(y_{k})$ from (3), and $a_{k+1}(a_{k+1}-1)\leq a_{k}^{2}$ for all $k\geq 0$ (which can be verified directly). Inequality (11), by rearrangement, shows that the function
@@
is a Lyapunov function in the sense that it is nonincreasing w.r.t. $k$. It follows that
@@
The proof of Theorem 6 is essentially the same as the original proof in [7]. We made the choice of step size more explicit, following [12]. Additionally, we reformulated the more general step size choice, $\beta_{k}$, considered in [12], to align with the format of the original proof in [7]. We believe this streamlined proof will be of interest to readers.

### v2

@@ 2 Preliminaries
The following well-known inequality for $L$-smooth functions will be used later.
@@
$f(y)$ $\leq f(x)+(-\alpha+\frac{L\alpha^{2}}{2})\left|\nabla f(x)\right|^{2},$ (3)
which, in particular, leads to
@@
when $\alpha=\frac{1}{L}$.
@@ 3 General convex functions
Let $p_{0}=0$. Then equation (7) also holds for $k=0$ by the initial condition $x_{0}=x_{-1}$ and $y_{0}=x_{0}$. By (7) and (5), it is straightforward to verify that
@@
$\,-2\alpha a_{k+1}(y_{k}-x^{*})\cdot\nabla f(y_{k})+a_{k+1}^{2}\alpha^{2}\left|\nabla f(y_{k})\right|^{2}.$ (8)
By (3), we have
$f(x_{k+1})\leq f(y_{k})+(-\alpha+\frac{L\alpha^{2}}{2})\left|\nabla f(y_{k})\right|^{2}.$ (9)
By convexity of $f$ (Definition 2), we have
$(y_{k}-x_{k})\cdot\nabla f(y_{k})\geq f(y_{k})-f(x_{k}).$ (10)
@@
$(y_{k}-x^{*})\cdot\nabla f(y_{k})\geq f(y_{k})-f^{*}.$ (11)
@@
$\quad-2\alpha a_{k+1}(f(y_{k})-f^{*})+a_{k+1}^{2}\alpha^{2}\left|\nabla f(y_{k})\right|^{2}$
$\leq\left|p_{k}+x_{k}-x^{*}\right|^{2}-2\alpha a_{k+1}^{2}(f(y_{k})-f^{*})$
$\quad+2\alpha a_{k}^{2}(f(x_{k})-f^{*})+a_{k+1}^{2}\alpha^{2}\left|\nabla f(y_{k})\right|^{2}$
$\leq\left|p_{k}+x_{k}-x^{*}\right|^{2}-2\alpha a_{k+1}^{2}(f(x_{k+1})-f^{*})$
$\quad+2\alpha a_{k}^{2}(f(x_{k})-f^{*})$
$\quad+a_{k+1}^{2}\alpha^{2}(-1+L\alpha)\left|\nabla f(y_{k})\right|^{2}$
@@
where the first inequality follows from substituting (10) and (11) into (8), the second inequality used $a_{k+1}(a_{k+1}-1)\leq a_{k}^{2}$ for all $k\geq 0$ (which can be verified directly), the third inequality follows from (9), and the fourth inequality holds because $\alpha\leq\frac{1}{L}$.
Inequality (12), by rearrangement, shows that
@@
is a Lyapunov function in the sense that it is nonincreasing with respect to $k$ along iterates of (5). It follows that
@@
The proof of Theorem 5 essentially follows the argument of the original proof in [7]. We made the choice of step size more explicit, following [12]. Additionally, we reformulated the more general step size choice, $\beta_{k}$, considered in [12], to align with the format of the original proof in [7]. We believe this streamlined proof will be of interest to readers.
@@ 5 Conclusions
{ack}
This work was partially supported by the NSERC of Canada and the CRC program. The author is grateful to Nicolas Boumal for identifying a typo in an earlier version that caused an issue in the proof of Theorem 5.